OpenStudy (anonymous):

A tank has 3 inlet pipes, each alone can fill the tank in 20 hours and 3 outlet pipes, each alone can empty the tank in 40 hours. At time t = 0, one inlet tap is opened. After 1 hour one outlet tap is also opened. The next hour, one more inlet and the next again outlet that means finally at the beginning of sixth hour all the 6 pipes are opened and no tap is closed after that until the tank is filled. How much time from the beginning is required for the tank to fill?

OpenStudy (anonymous):

1/20+1/40 -->first hour 6(3/40)=9/20

OpenStudy (shelby1290):

1ST PIPE WORKS X HR 2ND MAN (X-1) HR 3RD PIPE (X-2)HR AND SO ON TILL LAST HOUR. THEREFORE EQUTN BECOMES [ (X/20) ((X-2 )/20) ((X-4)/20) ) ]- [((X-1)/40 ((X-3)/40) ((X-5)/40))] = 1 X COMES 14 HRS 20 MIN

OpenStudy (shelby1290):

Your answer is 14 hours and 20 mins

OpenStudy (anonymous):

3q